# Toward Low-Latency, High-Fidelity Calibration of the LIGO Detectors with Enhanced Monitoring Tools

### v1

@@ 4.2 Real time Calibration Monitoring
The current observing run, O4, has brought the first ever deployment of a real-time monitoring system for calibrated LIGO data, CalMonitor. This development is one step along the road towards achieving low-latency, high-accuracy calibrated data that is robust to a changing interferometer. CalMonitor makes use of calibraiton lines injected by the Pcals. The frequencies corresponding to each of the calibration lines used for monitoring are known as monitoring line frequencies $f_{mon}$. At the monitoring line frequencies, the strain of the detector is dominated by the photon calibrator injection, $\tilde{h}_{\rm det}(f_{mon})\approx\tilde{h}_{\rm pcal}(f_{mon})$. The signal-to-noise ratio of these injected signals is such that a 256 second integration window yields an uncertainty in the injected signal coherence with the calibrated strain data of less than 2%. A real-time measure of the calibration systematic error $\eta_{R}$ at the monitoring line frequencies can then be constructed from the transfer function of the Pcal injected signal $h_{\rm pcal}$ in units of strain and the reconstructed detector strain data $h^{(\rm model)}_{\rm det}$ at each monitoring line frequency,
@@
Figure 2: Workflow for the data products associated with CalMonitor. The red containers indicate data streams that are required as inputs to CalMonitor. The yellow box represents all components of the DARM model needed for the calibration process. The light blue box CalUncertainty represents the process around constructing the calibration systematic error estimate. The black outlined box CalMonitor is the main software package used to compute calibration monitoring metrics. The remaining purple boxes are third-party software that is used to reduce, store, and display the calibration metrics.
CalMonitor uses the streaming software library gstlal to construct constant measurements of the calibration systematic error at the monitoring line frequencies [51]. In addition to calibration systematic error, we also construct and store metrics related to the state of the instrument, coherence of each transfer function measurement, values for the time-varying DARM model parameters [53, 54], and metrics related to the Pcal system, discussed more below. All of these constructed metrics are broadcast using the Kafka event streaming platform. The LVK software package ligo-scald is used to store these metrics from Kafka in InfluxDB [56]. We make use of the Grafana web application to visualize the real time calibration metrics. Figure 2 shows the data flow for the real time calibration monitoring system. In addition to visualizing calibration metrics using the Grafana web application, the transfer function measurements stored in InfluxDB are also pulled into our hourly calibration systematic error estimates, as will be discussed more in a forthcoming publication.
@@
Figure 4: Violin bode plots of the magnitude and phase of calibration systematic error as measured by the monitoring lines (see (11)) at the LIGO Hanford detector over the same time period as Figure 3. The top plots show the systematic error in the calibration when the detector was fully thermalized and no thermalization effects were measured in the detector sensing function. The bottom plots show the systematic error for a time when the detector was actively thermalizing. The larger spread in systematic error at the lower frequency monitoring lines is a result of the larger systematic error in the sensing function model during thermalization. The frequencies of the monitoring lines are shown along the top of each bode plot. The box on each violin plot extends from the first to third quartile. The yellow line shows the median value. The whiskers stretch from the box to the most distant data point that falls within 1.5 times the interquartile range of the box. Any points beyond the whiskers are considered outliers.
@@
Figs. 3 and 4 are representative of the types of plots displayed in real time on the Grafana web application. These plots display point estimates at specific frequencies for the systematic error of the calibrated data over the selected time period, as computed through (11), a violin plot representing the magnitude and phase of the systematic error of the calibration as a function of frequency over the selected time period, and time series of the time-varying DARM model parameters.
@@
The implementation of the real time calibration monitoring system has allowed for quick identification of problems with LIGO calibrated data and sanity checks on our computed calibration systematic error estimates. For example, in the early months of O4, the LIGO Hanford detector was running at higher power. The thermalization effects resulting from the higher laser power led to systematic errors in our modeled sensing function, as discussed in section 3. The real time calibration monitoring system clearly showed the trend of larger systematic errors at the beginning of lock stretches when the detector was thermalizing, as shown in Figs. 3 and 4. The effects are particularly apparent at the lowest frequency monitoring lines, where the thermalization effects have the largest impact on the detector sensing function.
@@
The real time calibraiton monitoring system has also been employed to study the Pcal absolute references. Using the characteristic feature of the interferometer that it cannot distinguish whether the X-end or the Y-end test mass was displaced, the two Pcal signals at each end station can be combined to provide a more accurate absolute calibration. This formalism was developed in O3 [29]. During O4, two frequencies, excited $0.1$ Hz apart, were continuously monitored at both the Hanford and Livingston sites. Fig 5 shows a time series of the quantity $\eta^{\rm X}_{R}/\eta^{\rm Y}_{R}$ for the LIGO Hanford detector, where
@@
where we assume $\tilde{h}^{(\rm model)}_{\rm det}(f_{\rm X})\approx\tilde{h}^{(\rm model)}_{\rm det}(f_{\rm Y})$ as $f_{\rm X}$ and $f_{\rm Y}$ are closely spaced. Continuous monitoring of $\eta^{\rm X}_{R}/\eta^{\rm Y}_{R}$ enabled the identification and correction of any relative error between the two Pcal calibrations at the end stations, ensuring accurate absolute calibration in real time.
Figure 5: Time series of the magnitude of the characteristic ratio $\eta^{\rm X}_{R}/\eta^{\rm Y}_{R}$ (see (13)) at the LIGO Hanford detector used to identify any relative errors between the two Pcal calibrations at each end station. This ratio should be one for identical Pcal calibrations. Fluctuations in this ratio overtime are caused primarily by noise in $h_{\rm det}^{\rm model}$. The data in this plot was collected using a configuration where $\eta^{\rm X}_{R}/\eta^{\rm Y}_{R}$ was computed using three 50% overlapping 240 second FFTs.
@@ 5 Calibration for the Next Generation of Gravitational-wave Detectors
The next-generation GW observatories, such as Cosmic Explorer [57] and Einstein Telescope [58], will extend GW observations to unprecedented distances, probing fundamental physics and astrophysical processes with unparalleled sensitivity. To fully realize their potential, robust and highly accurate calibration methods must be developed to ensure that raw interferometric data are precisely translated into astrophysical strain measurements.
Real-time calibration will be a crucial component of next-generation observatories. Future instruments will require real-time, high-accuracy calibration to enable immediate astrophysical analyses and better facilitate rapid responses for early warning alerts and multimessenger observations in a high-SNR regime. Achieving this requires advancements in detector calibration technology and methodology, including improved hardware, more sophisticated real-time monitoring, and efficient computational algorithms to interpret measurements and rapidly adjust the calibration.
@@
The development of these next-generation calibration methodologies builds upon the extensive progress made during the Advanced LIGO-Virgo-KAGRA observing runs. By refining systematic error modeling, improving real-time response characterization, and expanding the range of independent calibration techniques, the field is steadily advancing toward the goal of high-precision, real-time calibration. These innovations will be instrumental in unlocking the full scientific potential of Cosmic Explorer, Einstein Telescope, and other future gravitational-wave observatories, ensuring that they provide the most accurate and reliable data for astrophysical and fundamental physics discoveries.
@@ 6 Conclusion
Over the past decade, LIGO calibration has moved from producing calibrated strain data and the associated systematic error estimate on these data on a time scale of several months to a timescale of seconds. During the time of GW150914, we were able to achieve calibration errors less than $10\%$ in magnitude and $10$ deg in phase in the detection band. However, this required many months of work to produce recalibrated strain data from a well-understood DARM model. Currently, in O4, we are producing calibrated strain data with the same level of accuracy as was done for GW150914 but now with significantly lower latency. The low-latency strain data are available within seconds of data collection and preliminary uncertainty estimates are available on an hourly cadence. However, a thorough review of all data products is still a time-consuming process. The shift towards low-latency calibrated data that have a well-quantized systematic error estimate has eliminated the need for computationally expensive and time-consuming reruns of astrophysical analyses on recalibrated strain data that are delivered months after data collection.
The improved accuracy and latency of the calibrated strain data has been possible due to advances in the low-latency calibration pipeline, including improvements to the digital filters used in the calibration process and the subtraction of narrow-band noise in the calibrated strain data. Additionally, the development of a real-time monitoring system for the calibration has allowed for quick identification of problems in the calibration and provided real-time measurements of the calibration systematic error that inform the systematic error estimates. The real-time calibration monitoring system has also been deployed to measure any unexpected, systematic differences between the Pcals at the X- and Y-endstations.
@@ Appendix B The DARM Loop
The DARM loop, as introduced in section 2, is the feedback control loop that stabilizes the differential arm (DARM) degree of freedom in the interferometer. The response of the DARM loop, commonly called “the detector response”, must be characterized in order to achieve accurately calibrated differential arm length data from the interferometer. Ever since the km-scale LIGO detectors have begun operation, the parameters of the complex-valued, frequency-dependent detector response are observed to vary slowly over time and change the detector response for three primary reasons: (1) the alignment of the core optics (suspended from sophisticated multi-stage seismic isolation system in the Advanced LIGO era) drifts on the time-scale of seconds even while under servo control; (2) the core optics absorb a fraction of the circulating laser power in the Fabry-Pérot cavities (of order Watts) and thus distort under thermal load, changing the radii of curvature, and thus the detector response at various minute-long timescales; and (3) new for the advanced LIGO era, the alignment of the primary arm cavity optics are controlled using electrostatic actuators which are influenced by residual stray ions within the vacuum enclosure and accumulates over day-long timescales
@@ B.1 Sensing Function Model
where $\kappa_{C}$ is a slow time-varying, real-valued scale factor; $H_{C}$ is the overall scale of the sensing function; $f_{cc}$ is the so-called cavity pole frequency of the combined Fabry-Pérot and signal recycling cavities [61, 14], slowly varying in time; the second parenthetical term is a second-order phenomenological estimate of the measured response below 30 Hz [^3] This term encapsulates possible arm-to-signal-recycling cavity detuning coupled with non-ideal readout phase and is parameterized by $f_{s}$ and $Q$., where $f_{s}$ and $Q$ slowly vary in time; $C_{R}$ is any residual frequency response remaining in the true response of the signal processing electronics that cannot be compensated in the near real-time system; and $\tau_{C}$ accounts for the overall time delay. The time delay has three contributing components: a phase correction for approximating the full high-frequency response of the Fabry-Pérot cavity [62, 63] as a single pole [64], the light-travel time it takes to convey displacement of the end test masses down the arms, and any residual discrete delays incurred from sharing signals across the near-real-time control network of computers. The parameters in this model that are measured as a function of time are $\kappa_{C}$, $f_{cc}$, $f_{s}$, $Q$, and $\tau_{C}$.
@@ B.2 Actuation Function Model
Equation (15) becomes even more complex if several different quadruple suspension optics are used for actuation of the DARM loop. In principle, one could choose to construct a sophisticated actuation scheme that utilizes quadruple suspension Fabry-Pérot cavity optics of both input and end mirrors; different stages of different quadruple suspensions could be used or not; the relative strength on a given quadruple suspension stage could be increased or decreased digitally using the digital control system; and (adding to the complexity) there are four actuators per suspension stage, where different actuators could be manipulated separately using the digital system. At the time of GW150914, the DARM loop actuation was intentionally kept simple, but since then, the actuation scheme has slowly grown in complexity so that the actuation function model has also grown in complexity. A generalized, arbitrary actuation function model is beyond the scope of this article.
In [10], equation 4 used the symbol $\mathcal{K}_{i}$ instead of $H_{i}$ to represent the overall scale. Slow temporal variations $\kappa_{i}$ were left out of the definition of $A^{\textrm{(model)}}$ as they were not yet a part of the filters that produced $h_{\rm det}^{\textrm{(model)}}$ at the time. Their measured values were instead incorporated into the systematic error model, $\eta_{R}^{\textrm{(model)}}$. We imposed the expectation that $\kappa_{U}$ and $\kappa_{P}$ did not vary in time and could be lumped together as a common multiplicative factor, $\kappa_{PU}$. Finally, it was assumed that each stage had the same computational delay, $\tau_{i}=\tau_{A}$.

### v2

@@ 4.2 Real time Calibration Monitoring
The current observing run, O4, has brought the first deployment of a real-time monitoring system for calibrated LIGO data, CalMonitor. This development is one step along the road towards achieving low-latency, high-accuracy calibrated data that is robust to a changing interferometer. CalMonitor makes use of calibration lines injected by the Pcals. While Pcal calibration lines have been used throughout all of Advanced LIGO for various purposes, O4 marks the first time these lines have been analyzed in low-latency with results published to a live web application for real-time monitoring and review. Additionally, we are now injecting additional Pcal lines that are independent of the strain computation and serve only as monitors of calibration accuracy at specific frequencies. The frequencies corresponding to each of the calibration lines used for monitoring are known as monitoring line frequencies $f_{mon}$. At the monitoring line frequencies, the strain of the detector is dominated by the Pcal injection, $\tilde{h}_{\rm det}(f_{mon})\approx\tilde{h}_{\rm pcal}(f_{mon})$. The signal-to-noise ratio of these injected signals is such that a 256 second integration window yields an uncertainty in the injected signal coherence with the calibrated strain data of less than 2%. A real-time measure of the calibration systematic error $\eta_{R}$ at the monitoring line frequencies can then be constructed from the transfer function of the Pcal injected signal $h_{\rm pcal}$ in units of strain and the reconstructed detector strain data $h^{(\rm model)}_{\rm det}$ at each monitoring line frequency,
@@
Figure 3: Simplified workflow for the data products associated with CalMonitor. The red containers indicate timeseries data streams that are required as inputs to CalMonitor. The data products produced by CalMonitor are broadcast using the Kafka event streaming platform and then aggregated and stored in an InfluxDB backend using the ligo-scald software package. The details of these steps are omitted from this summary diagram, but the end result of calibration metric aggregation in a database is represented by the green cylinder. The Grafana web application is used to query InfluxDB and display calibration metrics in a real-time, user-friendly interface.
CalMonitor uses the streaming software library gstlal to perform real-time computations of the calibration systematic error at the monitoring line frequencies [55]. In addition to calibration systematic error, we also construct and store metrics related to the state of the instrument, coherence of each transfer function measurement, values for the time-varying DARM model parameters [58, 59], and metrics related to the Pcal system, discussed more below. All of these constructed metrics are broadcast using the Kafka event streaming platform. The LVK software package ligo-scald is used to store these metrics from Kafka in InfluxDB [61]. We make use of the Grafana web application to visualize the real-time calibration metrics. Figure 3 shows the data flow for the real time calibration monitoring system. In addition to visualizing calibration metrics using the Grafana web application, the transfer function measurements stored in InfluxDB are also pulled into our hourly calibration systematic error estimates, as will be discussed more in a forthcoming publication.
@@
Figure 5: Violin bode plots of the magnitude and phase of calibration systematic error as measured by the monitoring lines (see (11)) at the LIGO Hanford detector over the same time period as figure 4. The top plots show the systematic error in the calibration when the detector was fully thermalized and no thermalization effects were measured in the detector sensing function. The bottom plots show the systematic error for a time when the detector was actively thermalizing. The larger spread in systematic error at the lower frequency monitoring lines is a result of the larger systematic error in the sensing function model during thermalization. The frequencies of the monitoring lines are shown along the top of each bode plot. The box on each violin plot extends from the first to third quartile. The black line shows the median value. The whiskers stretch from the box to the most distant data point that falls within 1.5 times the interquartile range of the box. Any points beyond the whiskers are considered outliers.
@@
Figures 4 and 5 are representative of the types of plots displayed in real time on the Grafana web application. These plots display point estimates at specific frequencies for the systematic error of the calibrated data over the selected time period, as computed through (11), a violin plot representing the magnitude and phase of the systematic error of the calibration as a function of frequency over the selected time period, and time series of the time-varying DARM model parameters.
@@
The implementation of the real-time calibration monitoring system has allowed for quick identification of problems with LIGO calibrated data and sanity checks on our computed calibration systematic error estimates. For example, in the early months of O4, the LIGO Hanford detector was operating with a higher laser power. The thermalization effects resulting from the higher laser power led to systematic errors in our modeled sensing function, as discussed in section 3. The real-time calibration monitoring system clearly showed the trend of larger systematic errors during times of thermalization, as shown in figures 4 and 5. The effects are particularly apparent at the lowest frequency monitoring lines, where the thermalization effects have the largest impact on the detector sensing function.
Figure 6: Time series of the magnitude of the characteristic ratio $\eta^{\rm X}_{R}/\eta^{\rm Y}_{R}$ (see (13)) at the LIGO Hanford detector used to identify any relative errors between the two Pcal calibrations at each end station. This ratio should be one for identical Pcal calibrations. The deviation from unity is due to uncertainties in the contributing factors in Pcal calibrations, not common to the two end stations. The data in this plot was collected using a configuration where $\eta^{\rm X}_{R}/\eta^{\rm Y}_{R}$ was computed using three 50% overlapping 240 second FFTs. The mean of the variations in the ratio is $\sim 1.0005$, indicating that there is no significant relative Pcal errors.
@@
The real-time calibration monitoring system has also been employed to study the Pcal absolute references. Using the characteristic feature of the interferometer that it cannot distinguish whether the X-end or the Y-end test mass was displaced, the two Pcal signals at each end station can be combined to provide a more accurate absolute calibration. This formalism was developed in O3 [36]. During O4, two frequencies $f_{\rm X}$ and $f_{\rm Y}$, excited $0.1$ Hz apart, were continuously monitored at both the Hanford and Livingston sites. Figure 6 shows a time series of the quantity $\eta^{\rm X}_{R}/\eta^{\rm Y}_{R}$ for the LIGO Hanford detector, where
@@
where we assume $\tilde{h}^{(\rm model)}_{\rm det}(f_{\rm X})\approx\tilde{h}^{(\rm model)}_{\rm det}(f_{\rm Y})$ as $f_{\rm X}$ and $f_{\rm Y}$ are closely spaced. Continuous monitoring of $\eta^{\rm X}_{R}/\eta^{\rm Y}_{R}$ enabled the identification and correction of any relative error between the two Pcal calibrations at the end stations, ensuring an accurate absolute calibration in real time.
The real-time calibration monitoring system became a critical tool for the LIGO calibration team during the O4 observing run. The infrastructure for CalMonitor was designed to be flexible enough to allow for expanded calibration metrics and monitoring in future observing runs. Similar systems will be required as we move into the next-generation of GW detectors as well.
@@ 5 Calibration for the Next Generation of Gravitational-wave Detectors
The next-generation GW observatories, such as Cosmic Explorer [62] and Einstein Telescope [63], will extend GW observations to unprecedented distances, probing fundamental physics and astrophysical processes with unparalleled sensitivity. To fully realize their potential, robust and highly accurate calibration methods must be developed to ensure that raw interferometric data are precisely translated into astrophysical strain measurements [30, 64].
Real-time calibration will be a crucial component of next-generation observatories. Future instruments will require real-time, high-accuracy calibration to enable immediate astrophysical analyses and better facilitate rapid responses for early-warning alerts, which are alerts that are sent out before the merger phase of the binary coalescence, and multimessenger observations in a high signal-to-noise ratio regime [65]. Achieving this requires advancements in detector calibration technology and methodology, including improved hardware, more sophisticated real-time monitoring, and efficient computational algorithms to interpret measurements and rapidly adjust the calibration. Alternative calibration schemes, such as astrophysical calibration, also provide promising methods for calibrating the next generation detectors [66, 67].
@@
The development of these next-generation calibration methodologies builds upon the extensive progress made during the Advanced LIGO-Virgo-KAGRA observing runs. By refining systematic error modeling, improving real-time response characterization, and expanding the range of independent calibration techniques, the field is steadily advancing toward the goal of high-precision, real-time calibration. These innovations will be instrumental in unlocking the full scientific potential of Cosmic Explorer, Einstein Telescope, and other future GW observatories, ensuring that they provide the most accurate and reliable data for astrophysical and fundamental physics discoveries.
@@ 6 Conclusion
Over the past decade, LIGO calibration has moved from producing calibrated strain data and the associated systematic error estimate on these data on a time scale of several months to a timescale of seconds for the data and hours for the systematic error estimate. During the time of GW150914, we were able to achieve calibration errors less than $10\%$ in magnitude and $10$ deg in phase in the detection band. However, this required many months of work to produce recalibrated strain data from a well-understood DARM model. In O4, we are producing calibrated strain data with the same level of accuracy as was done for GW150914 but now with significantly lower latency. The low-latency strain data are available within seconds of data collection and preliminary uncertainty estimates are available on an hourly cadence. However, a thorough review of all data products is still a time-consuming process. The shift towards low-latency calibrated data that have a well-quantized systematic error estimate has eliminated the need for computationally expensive and time-consuming reruns of astrophysical analyses on recalibrated strain data that are delivered months after data collection.
The improved accuracy and latency of the calibrated strain data has been possible due to advances in the low-latency calibration pipeline, including improvements to the digital filters used in the calibration process and the subtraction of narrow-band noise in the calibrated strain data. Additionally, the development of a real-time monitoring system for the calibration has allowed for quick identification of problems in the calibration and provided real-time measurements of the calibration systematic error that inform the systematic error estimates. The real-time calibration monitoring system has also been deployed to measure any unexpected, systematic differences between the Pcals at the X- and Y-end stations. The real-time calibration monitoring system will be expanded to include new calibration metrics, such as metrics related to line subtraction and thermalization effects.
@@ Appendix B The DARM Loop
The DARM loop, as introduced in section 2, is the feedback control loop that stabilizes the differential arm (DARM) degree of freedom in the interferometer. The response of the DARM loop, commonly called “the detector response”, must be characterized in order to achieve accurately calibrated differential arm length data from the interferometer. Ever since the km-scale LIGO detectors have begun operation, the parameters of the complex-valued, frequency-dependent detector response are observed to vary slowly over time and change the detector response for three primary reasons: (1) the alignment of the core optics (suspended from sophisticated multi-stage seismic isolation system in the Advanced LIGO era) drifts on the time-scale of seconds even while under servo control; (2) the core optics absorb a fraction of the circulating laser power in the Fabry-Perot cavities (of order watts) and thus distort under thermal load, changing the radii of curvature, and thus the detector response at various minute-long timescales; and (3) new for the advanced LIGO era, the alignment of the primary arm cavity optics are controlled using electrostatic actuators which are influenced by residual stray ions within the vacuum enclosure and accumulates over day-long timescales
@@ B.1 Sensing Function Model
where $\kappa_{C}$ is a slow time-varying, real-valued scale factor; $H_{C}$ is the overall scale of the sensing function; $f_{cc}$ is the so-called cavity pole frequency of the combined Fabry-Perot and signal recycling cavities [70, 14], slowly varying in time; the second parenthetical term is a second-order phenomenological estimate of the measured response below 30 Hz [^3] This term encapsulates possible arm-to-signal-recycling cavity detuning coupled with non-ideal readout phase and is parameterized by $f_{s}$ and $Q$., where $f_{s}$ and $Q$ slowly vary in time; $C_{R}$ is any residual frequency response remaining in the true response of the signal processing electronics that cannot be compensated in the near real-time system; and $\tau_{C}$ accounts for the overall time delay. The time delay has three contributing components: a phase correction for approximating the full high-frequency response of the Fabry-Perot cavity [71, 72] as a single pole [73], the light-travel time it takes to convey displacement of the end test masses down the arms, and any residual discrete delays incurred from sharing signals across the near-real-time control network of computers. The parameters in this model that are measured as a function of time are $\kappa_{C}$, $f_{cc}$, $f_{s}$, $Q$, and $\tau_{C}$.
@@ B.2 Actuation Function Model
Equation (15) becomes even more complex if several different quadruple suspension optics are used for actuation of the DARM loop. In principle, one could choose to construct a sophisticated actuation scheme that utilizes quadruple suspension Fabry-Perot cavity optics of both input and end mirrors; different stages of different quadruple suspensions could be used or not; the relative strength on a given quadruple suspension stage could be increased or decreased digitally using the digital control system; and (adding to the complexity) there are four actuators per suspension stage, where different actuators could be manipulated separately using the digital system. At the time of GW150914, the DARM loop actuation was intentionally kept simple, but since then, the actuation scheme has slowly grown in complexity so that the actuation function model has also grown in complexity. A generalized, arbitrary actuation function model is beyond the scope of this article.
In [10], equation (4) used the symbol $\mathcal{K}_{i}$ instead of $H_{i}$ to represent the overall scale. Slow temporal variations $\kappa_{i}$ were left out of the definition of $A^{\textrm{(model)}}$ as they were not yet a part of the filters that produced $h_{\rm det}^{\textrm{(model)}}$ at the time. Their measured values were instead incorporated into the systematic error model, $\eta_{R}^{\textrm{(model)}}$. We imposed the expectation that $\kappa_{U}$ and $\kappa_{P}$ did not vary in time and could be lumped together as a common multiplicative factor, $\kappa_{PU}$. Finally, it was assumed that each stage had the same computational delay, $\tau_{i}=\tau_{A}$.